OpenStudy (anonymous):

The conical hat shown here is a traditional sun-protective straw hat originating in eastern Asia. What is the approximate outer surface area of one of these hats that has a radius of 9 inches and a slant height of 12 inches? Keep in mind that there is an opening in the hat where the base surface would normally be. Show the steps of your solution and explain your work.

OpenStudy (anonymous):

Don't know if that's important but here you go:

OpenStudy (anonymous):

HELP PLEASE.. O_o

OpenStudy (anonymous):

what the heck is the picture for?

OpenStudy (anonymous):

Idk, it came with the problem, just answer my question. Dx

OpenStudy (anonymous):

im bad at math:(

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

..Eh.. Wow, then why'd you write on here?

OpenStudy (anonymous):

to make you feel better

OpenStudy (anonymous):

The surface area of a cone (excluding the base) is pi*r*s, where r = radius and s = slant height So in this case, we have pi*9*12 = 96*pi =~ 301.59 in^2